OpenStudy (anonymous):

A certain rectangular prism has a height of h inches. It has a width that is 3 less than its height and a length that is 4 inches more than its height. If the area of the base is 60 in^2, what is the volume in cubic inches. (Answer is 480, but I don't understand the process)

OpenStudy (anonymous):

Willing to give a medal or whatever if someone tells me how to!

OpenStudy (anonymous):

Help please

OpenStudy (hurgleman):

You know that the base is 60 60 = l * w you know that width is height - 3 inches w = h-3 you know that length is h+4 inches l=h+4 So now you plug in length and width with those equations. 60 = (h+4) * (h-3) Foil the expressions. 60 = h^2 + h -12 expressing h as x 60=x^2+x-12 or 72=h^2+h Set one side to 0 0 = h^2 + h - 72 Then, you can use the quadratic formula to identify what 2 factors this breaks down to. \[(-1 +\sqrt{1-(4*-72})/2\] so (-1+- 17)/2 your answers are -9 or 8. h = 8 or h=-9 Since it's volume, you can't use a negative number, so h=8

OpenStudy (hurgleman):

From there you can plug in 8 to anywhere that H is. V = 8 * (8-3) * (8+4) = 480 480 = 8 * 5 * 12